OpenStudy (anonymous):

Graph the parametrized curve given by x=t-2, y=-2+square root(5t), 0 is less than or equal to t which is less than or equal to 4. Indicate the direction in which the curve is traced.

OpenStudy (anonymous):

@abb0t please help

OpenStudy (abb0t):

Well, what you can do is eliminate \(\sf \color{green}{t}\) in one of them to substitute for another.

OpenStudy (abb0t):

So, \(\sf \color{red}{t=x+2}\) plug that into \(\sf \color{blue}{y=-2+ \sqrt{5(x+2)}}\)

OpenStudy (anonymous):

y=-2+ square root of (5x -10)

OpenStudy (abb0t):

Actually, it might be easier to see your graph if you elimintet the bottom function. Lol sorry.

OpenStudy (abb0t):

It should be a parabola.

OpenStudy (anonymous):

i dont think its a parabola because of the square root

OpenStudy (abb0t):

well, if you solve for \(t\) in the y, and plug it into the top one x = t-2, you should get a parabola.

OpenStudy (anonymous):

Would it be t=(y^2 + 4y + 4)/5

OpenStudy (anonymous):

@satellite73 please help

OpenStudy (abb0t):

Which is a parabola. Yes.

OpenStudy (anonymous):

where do i plug in the x=t-2?